# Effects of anisotropy in an anisotropic extension of $w$CDM model

### v1

@@ V Concluding Remarks
In the present work, we have derived CMB independent observational constraints on two simplest extensions of $\Lambda$CDM model, namely $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models from recent data sets including BAO, CC, BBN, PP and PPSH0ES in two combinations: BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES. We have obtained that DE EoS parameter favours quintessence form of DE for $w$CDM+$\Omega_{\rm\sigma 0}$ model from both the data combinations at $68\%$ CL as well as $95\%$ CL. For the $w$CDM model, the mean value of $w_{\rm de0}$ at $68\%$ CL lies in quintessence region with BAO+CC+BBN+PP data combination whereas, it touches the phantom barrier at $95\%$ CL with the same data combination. Again, for the same model with BAO+CC+BBN+PPSH0ES data combination, the mean value of $w_{\rm de0}$ lies in phantom region $(w_{\rm de0}<-1)$ at $95\%$ CL. Also, we notice a strong positive correlation between $w_{\rm de0}$ and anisotropy parameter ($\Omega_{\rm\sigma 0}$) with both data combinations, see Figure 1. The upper bounds on $\Omega_{\rm\sigma 0}$ are of the order $10^{-13}$ from both the data combinations. The constant EoS of DE significantly affects present day matter density parameter and provides lower values for both the models (as compared to models investigated in [60] with cosmological constant DE) from both the data combinations.
We have obtained higher values of Hubble constant with $H_{\rm 0}=71.51\pm 0.83$ and $H_{\rm 0}=72.24\pm 0.82$ $\rm km\,s^{-1}\,Mpc^{-1}$ at 68% CL for $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models, respectively with BAO+CC+BBN+PPSH0ES data. The obtained value of $H_{0}$ in both the considered models from BAO+CC+BBN+PPSH0ES are consistent with SH0ES measurements ($H_{0}^{\rm R22}$), and thus $H_{0}$ tension is relieved. Further, we notice that inclusion of anisotropy parameter in the $w$CDM model reduces the $H_{0}$ tension by $\sim 2\sigma$ and $\sim 1\sigma$ with BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES data combinations, respectively. Overall, we have shown that how the anisotropic extensions of standard cosmological model can influence the estimation of $H_{0}$, and contribute significantly to the resolution of the $H_{0}$ tension.

### v2

@@ V Concluding Remarks
In the present work, we have derived CMB independent constraints on two simplest extensions of the $\Lambda$CDM model, namely the $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models from recent data sets including BAO, CC, BBN, PP and PPSH0ES in two combinations: $D_{1}$ and $D_{2}$. We have obtained that the DE EoS parameter favors the quintessence form for the $w$CDM+$\Omega_{\rm\sigma 0}$ model from both data combinations at $68\%$ CL as well as $95\%$ CL. For the $w$CDM model, the mean value of $w_{\rm de0}$ at $68\%$ CL lies in the quintessence region whereas, it touches the phantom barrier at $95\%$ CL with $D_{1}$. Again, for the same model with $D_{2}$, the mean value of $w_{\rm de0}$ lies in the phantom region $(w_{\rm de0}<-1)$ at $95\%$ CL. Also, we notice a strong positive correlation between $w_{\rm de0}$ and the anisotropy parameter ($\Omega_{\rm\sigma 0}$) in both cases, see Figure 1. The upper bounds on $\Omega_{\rm\sigma 0}$ are of the order $10^{-13}$ from $D_{1}$ and $D_{2}$. The constant EoS of the DE significantly affects the present matter density parameter and provides lower values for both models (as compared to models investigated in [60] with a cosmological constant) from both data.
We have obtained higher values of the Hubble constant with $H_{\rm 0}=71.51\pm 0.83$ and $H_{\rm 0}=72.24\pm 0.82$ $\rm km\,s^{-1}\,Mpc^{-1}$ at 68% CL for the $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models, respectively with $D_{2}$. The obtained value of $H_{0}$ in both the considered models from $D_{2}$ are consistent with SH0ES measurements ($H_{0}^{\rm R22}$), and thus the $H_{0}$ tension is relieved. Further, we notice that the inclusion of the anisotropy parameter in the $w$CDM model reduces the $H_{0}$ tension by $\sim 2\sigma$ and $\sim 1\sigma$ with $D_{1}$ and $D_{2}$, respectively. Overall, we have shown how the anisotropic extensions of the standard cosmological model can influence the estimation of $H_{0}$, and contribute significantly to the resolution of the $H_{0}$ tension.
Data Availability Statement
No data is associated with this manuscript.
Acknowledgements.
The authors gratefully thank the reviewers for constructive comments and fruitful suggestions to enhance the quality of the manuscript.